# Hound: Locating Cryptographic Primitives in Desynchronized Side-Channel Traces Using Deep-Learning

### v1

@@ Abstract
Side-channel attacks allow to extract sensitive information from cryptographic primitives by correlating the partially known computed data and the measured side-channel signal. Starting from the raw side-channel trace, the preprocessing of the side-channel trace to pinpoint the time at which each cryptographic primitive is executed, and, then, to re-align all the collected data to this specific time represent a critical step to setup a successful side-channel attack. The use of hiding techniques has been widely adopted as a low-cost solution to hinder the preprocessing of side-channel traces thus limiting side-channel attacks in real scenarios. This work introduces Hound, a novel deep learning-based pipeline to locate the execution of cryptographic primitives within the side-channel trace even in the presence of trace deformations introduced by the use of dynamic frequency scaling actuators. Hound has been validated through successful attacks on various cryptographic primitives executed on an FPGA-based system-on-chip incorporating a RISC-V CPU, while dynamic frequency scaling is active. Experimental results demonstrate the possibility of identifying the cryptographic primitives in DFS-deformed side-channel traces.
@@
Side-channel analysis, dynamic frequency scaling, deep-learning, locating of cryptographic operations.
@@ I Introduction
Side-channel analysis (SCA) poses a significant threat to modern cryptographic implementations, even when the underlying algorithms boast mathematical soundness. By exploiting vulnerabilities in the physical implementation of the cryptographic primitives (CPs), side-channel attacks can exploit the unintended information leakage generated from electronic devices during the execution of CPs [1]. The leakage, often manifested in power consumption [2] or electromagnetic emissions [3], can be analyzed by attackers to potentially reveal confidential information by correlating it with partially known processed data.
Traditionally, properly implemented masking [4, 5] and hiding [6, 7, 8] countermeasures offer robust protection against side-channel attacks, leading to successful security evaluations. Masking countermeasures split sensitive intermediate values into independent shares that are processed separately to significantly reduce the correleation between each share and the secret key. Hiding countermeasures introduce randomness into the targeted side-channel signal to degrade the signal-to-noise ratio.
Over the past two decades, researchers have proposed various methods to enhance the effectiveness of side-channel analysis, such as differential power analysis (DPA) [2], template attacks (TA) [9], correlation power analysis (CPA) [10], as well as DL-based solutions [11, 12]. Notably, all the proposed techniques share two key requirements. First, they require a large number of executions of the same CP with different inputs. Second, the attacker needs to locate and align in time all the CP executions in the side-channel trace to feed the attack method of choice.
@@
In contrast, attacking the implementation of a CP in the real-world presents a more significant challenge. Indeed, the attacker needs to locate the CP within the side-channel trace. Specific settings still permit a rough alignment of the measured side-channel signal with CP execution even in real world scenarios, e.g., by leveraging specific logic events happening in the computing platform or by using pattern-matching techniques applied to the side-channel trace to generate so-called virtual triggers [15, 16].
However, the importance of pinpointing the executions of the CPs in the side-channel trace even without a triggering infrastructure motivated several contributions, i.e., [17, 18]. In a similar manner, the continuous improvement in the attack techniques, fueled the use of hiding techniques to hinder the task of locating a CP, thus preventing the localization of the CPs in the side-channel trace as an additional security countermeasure.
[19] is the seminal work that presented a deep-learning pipeline to locate the CPs within side-channel traces deformed by means of random delay coutermeasures. However, random delay introduces a limited trace deformation and its implementation requires custom hardware or software implementations inducing non-negligible performance overheads [20]. In contrast, Dynamic Frequency Scaling (DFS) actuators offer a cheap and standard solution that has been widely investigated to implement hiding techniques. The possibility of employing hundrends of operating frequencies allows to effectively deform the side-channel trace at the point of preventing the correct pinpointing of the CP and, consequently, invalidating the possibility of attacking the target device even if the attacker can deploy an effective side-channel attack.
@@
  • We evaluate the effectiveness of our solution across various cryptographic primitives. Furthermore, we perform successful side-channel attacks after CP localization and alignment to validate the quality of the results also comparing our solution against state-of-the-art techniques.
  • To facilitate further research and ensure reproducibility, we have released Hound as an open-source tool along with a dataset of relevant side-channel traces. The tool and dataset are available at 111link details withheld for anonymity.
@@ II Related Works
Figure 2: Focus on the proposed Hound training pipeline, divided into Dataset Building and CNN Training. $Class_{0}$, $class_{1}$ and, $class_{2}$ contain CP start part windows, CP spare parts windows and, noise windows respectively.
However, using architectural-level techniques to morph the power trace represents an easy-to-implement and effective countermeasure to deceive pattern-matching-based solutions. For instance, employing time-sharing multithreading on a single-core microcontroller [22] can obstruct an attacker’s ability to accurately locate CPs, with interrupt service routines further modifying the side-channel trace’s form. Consequently, more sophisticated methods for detecting CPs in the side-channel trace have been developed. [19] developed a deep-learning approach for identifying deformed CPs through random delay, employing a Convolutional Neural Network (CNN) trained to differentiate CPs from the remaining portions of the power trace. Even though random delay-based hiding techniques generate some degree of CP deformation, they are limited in their ability to introduce significant variability in the power trace [20].
Despite these advancements, current state-of-the-art techniques struggle with highly deformed traces, particularly when the computing platform employs effective hiding countermeasures like DFS [23]. These countermeasures profoundly alter the timing of operations and continuously morph the side-channel signal, making it challenging to isolate CPs using traditional methods. To address this limitation, we propose a deep-learning-based technique for CP location in power traces collected from platforms implementing DFS as a countermeasure.
@@ III Methodology
Figure 1 depicts the proposed Hound methodology consisting of a training and inference pipeline. Figure 1a reports the Training Pipeline which aims at creating a classifier that can sort a window of the side-channel trace into three categories: start of a CP, spare part of a CP, or just noise (see Figure 2). This framework locates where the CPs begin in the side-channel trace without removing the obfuscation countermeasure. We assume that the attacker leverages an exact copy of the target device to collect a noise trace and several cipher traces for the training. The noise trace is obtained by running different applications that are not the CP. Such information is necessary to train a neural network to distinguish the beginning of the CPs from the execution of other applications. Each cipher trace is gathered while a single CP is running, where the attacker decides what data and keys to use. All measurements are collected with the obfuscation mechanism active since the attacker cannot turn it off. Starting with measured raw traces, the Dataset Building step creates a database of windows from the noise and cipher traces, each window being $N$ samples long. This database is then used to train a CNN classifier.
Figure 1b reports the Inference Pipeline. At inference time, the goal is to identify the CPs within a novel side-channel trace from the target device by employing the trained CNN. This pipeline is divided into three key phases: Sliding Window Classification, Screening, and Alignment. The inference pipeline takes in a single side-channel trace and segments it to mark the start of each CP. Initially, the Slicing stage processes a side-channel trace and generates a series of windows to classify with the trained CNN. Starting from the classified windows, representing a segmentation of the input trace, the Screening phase produces a sequence of time instants marking the start of each CP in the trace. Finally, the Alignment phase chunks the original side-channel trace based on these outputs, arranging the identified CPs accordingly.
The details related to the creation of the dataset and the CNN architecture are discussed in Section III-A, while the sliding window classification, and the screening procedure are detailed in Section III-B.
@@ III-A Training Pipeline
Dataset Building - The first stage in the training pipeline is Dataset Building, which receives as input a collection of side-channel traces and generates a dataset suitable for training the CNN. The attacker uses a clone of the target computing platform to create a noise trace and a set of cipher traces. The noise trace captures the execution of various applications, excluding the target CP. Each cipher trace is collected while executing a single CP, where the attacker controls the plaintext and secret key. Notably, the attacker cannot disable the obfuscation mechanism, so the DFS remains active during all traces collection. Considering the proposed security assumptions, the attacker can only execute a chosen application and measure the corresponding side channel on the clone device. However, the attacker can probe the clone device to detect the start of each CP. In contrast, probing is not required for the noise trace. Once trained, the CNN works on the target architecture that implements DFS without the need for probing. For each cipher trace $i$ of length $L_{i}$ samples, the starting $N$ samples are labeled as start of the CP (see $class_{0}$ in Figure 2). The remaining $L_{i}-N$ samples are equally split into consecutive windows of width $N$ and labeled as spare part of the CP (see $class_{1}$ in Figure 2). Moreover, we extract a random set of $N$-sample windows from the noise trace and we label each of them as noise (see $class_{2}$ in Figure 2). Although the methodology aim to identify only the beginning of each CP, the division into three classes allows CNN to maximize the distinction over the three cases.
Convolutional Neural Network - The architecture of the 1D CNN is adapted from the CNN proposed in [19], available open-source.
The CNN takes a window of $N$ samples from the side-channel trace as input and outputs a classification score vector. The CNN’s structure is organized into six sequential components: a convolutional block, a pair of residual blocks, a global average pooling layer, a fully-connected block, and a softmax layer. Each convolutional block features a 1D convolutional layer, a batch normalization layer [24], and a ReLU activation function [relu]. The residual blocks [25], each containing two convolutional blocks, are augmented with shortcut connections to perform element-wise feature summation. All the convolutional layers implement a kernel size of 64, a stride of 1, and zero padding to maintain the sample count to $N$. The first convolutional layer and the one in the first residual block implement 16 filters, while the subsequent residual block doubles the filter count to 32. The global average pooling layer averages the obtained features over the temporal dimension $N$. The feature vector is then processed through two fully-connected layers with a ReLU activation function, which outputs a linear vector. The softmax layer completes the architecture, generating a $3$-dimensional vector with the classification scores for each class.
@@ III-B Inference Pipeline
The Inference Pipeline, detailed in Figure 3, is divided into Sliding Window Classification, Screening, and Alignment stages. Leveraging the trained CNN classifier, the pipeline locates the starting sample of each CP present in a new side-channel trace. Notably, the pipeline is designed to work with highly obfuscated side-channel signals via DFS.
Sliding Window Classification - The first stage in the inference pipeline is Sliding Window Classification. It takes a new side-channel trace, slices it into $N$-sample windows, and uses the trained CNN to output a classification score to label each window as start of the CP, spare part of the CP, or noise (see Sliding Window Classification in Figure 3). The Slicing block implements a sliding window procedure to slice the side-channel trace in input. It takes three inputs, i.e., the side-channel trace, the size of the sliding window ($N$), and the stride ($s$), and outputs an ordered set of $N$-sample windows to feed the CNN.
@@
Screening - The Screening stage refines the segmentation output from the sliding window classification block to identify the beginning of the CPs in the side-channel trace. Algorithm 1 highlights the screening procedure. The screening algorithm (see Algorithm 1) takes the segmentation output from the sliding window classification block, denoted as $seg$, and some parameters, i.e., the stride $s$ used for the sliding window, the kernel size $k$ for the polish procedure (see line 10 of Algorithm 1), and the average CP length $avgCP$. In the end, it returns the list $starts$ of samples corresponding to the beginning of each CP instance in the processed side-channel trace.
The algorithm is an iterative process that involves three main steps: POLISH, EXTRACT, and REFINE. The core idea behind Algorithm 1 is to start with an aggressive polish and incrementally refine it by attempting to uncover new CPs. The polishing step (see line 10 of Algorithm 1) employs a majority filter across the whole segmentation. This filter traverses the input signal for a given kernel size $k$, replacing each point with the most frequently occurring value within a $k$-sized window. The extraction step (see line 14 of Algorithm 1) identifies the initial sample of each CP. It pinpoints the indices of the falling edge down to class 0, meaning those samples with a value of 0 where the preceding value is not 0. These samples are then scaled by $s$, i.e., the stride in the sliding window classification. Finally, the refining step (see line 23 of Algorithm 1) fine-tunes the algorithm’s parameters. It reduces kernel size $k$ and generates a list of sub-segmentations for further iteration. The list is given by those parts of the input segmentation for which there might still be two or more hidden CPs. The process begins by determining the minimum length of a CP $minCP$ (see line 25 of Algorithm 1), which is derived by comparing the average length of a CP provided as input with the starting points of CPs identified thus far. Next, the algorithm examines pairs of consecutive CP starting points (see line 28 of Algorithm 1). If the distance between them exceeds twice the $minCP$, it suggests the likelihood of additional CPs being present in such an interval. Consequently, the corresponding sub-segmentation is added to the list for further processing.
@@ IV Experimental Evaluation
TABLE I: Parameters for each pipeline stage over all the tested cipher.
Cipher Noise Parameters Metrics
Applications avgCP k N s Hits Mean IoU Std IoU
AES 145k 150 10k 62 100% 97.01% 1.65%
@@
This section presents the experimental results of Hound as organized in four parts. Section IV-A outlines the experimental setup. Experimental results targeting the CNN training and inference pipelines are discussed in Section IV-B and Section IV-C, respectively. A comprehensive example workflow, including a successful side-channel attack and a comparison with state-of-the-art proposals, is provided in Section IV-D.

### v3

@@ Abstract
Side-channel attacks allow the extraction of sensitive information from cryptographic primitives by correlating the partially known computed data and the measured side-channel signal. Starting from the raw side-channel trace, the preprocessing of the side-channel trace to pinpoint the time at which each cryptographic primitive is executed, and, then, to re-align all the collected data to this specific time represent a critical step to setup a successful side-channel attack. The use of hiding techniques has been widely adopted as a low-cost solution to hinder the preprocessing of side-channel traces, thus limiting side-channel attacks in real scenarios. This work introduces Hound, a novel deep-learning-based pipeline to locate the execution of cryptographic primitives within the side-channel trace even in the presence of trace deformations introduced by the use of dynamic frequency scaling actuators. Hound has been validated through successful attacks on various cryptographic primitives executed on an FPGA-based system-on-chip incorporating a RISC-V CPU while dynamic frequency scaling is active. Experimental results demonstrate the possibility of identifying the cryptographic primitives in DFS-deformed side-channel traces.
@@
Side-channel analysis, dynamic frequency scaling, deep-learning, locating of cryptographic primitives.
@@ I Introduction
Side-channel attacks pose a significant threat to modern cryptographic implementations, even when the underlying algorithms boast mathematical soundness. By exploiting vulnerabilities in the physical implementation of the cryptographic primitives (CPs), side-channel attacks can exploit the unintended information leakage generated from electronic devices during the execution of CPs [1]. The leakage, often manifested in power consumption [2] or electromagnetic emissions [3], can be analyzed by attackers to potentially reveal confidential information by correlating it with partially known processed data.
Traditionally, properly implemented masking [4, 5] and hiding [6, 7, 8] countermeasures offer robust protection against side-channel attacks, leading to successful security evaluations. Masking countermeasures split sensitive intermediate values into independent shares that are processed separately to reduce significantly the correlation between each share and the secret key. Hiding countermeasures introduce randomness into the targeted side-channel signal to degrade the signal-to-noise ratio.
Over the past two decades, researchers have proposed various methods to enhance the effectiveness of side-channel analysis, such as differential power analysis (DPA) [2], template attacks [9], correlation power analysis (CPA) [10], as well as deep-learning-based solutions [11, 12]. Notably, all the proposed techniques share two key requirements. First, they require a large number of executions of the same CP with different inputs. Second, the attacker needs to locate and align in time all the CP executions in the side-channel trace to feed the attack method of choice.
@@
In contrast, attacking the implementation of a CP in the real world presents a more significant challenge. Indeed, the attacker needs to locate the CP within the side-channel trace. Specific settings still permit a rough alignment of the measured side-channel signal with CP execution even in real-world scenarios, e.g., by leveraging specific logic events happening in the computing platform or by using pattern-matching techniques applied to the side-channel trace to generate so-called virtual triggers [15, 16].
However, the importance of pinpointing the executions of the CPs in the side-channel trace even without a triggering infrastructure motivated several contributions, i.e., [17, 18]. In a similar manner, the continuous improvement in the attack techniques fueled the use of hiding techniques to hinder the task of locating a CP, thus preventing the localization of the CPs in the side-channel trace as an additional security countermeasure.
[19] is the seminal work that presented a deep-learning pipeline to locate the CPs within side-channel traces deformed by means of random delay countermeasures. However, random delay introduces a limited trace deformation, and its implementation requires custom hardware or software implementations inducing non-negligible performance overheads [20]. In contrast, Dynamic Frequency Scaling (DFS) actuators offer a cheap and standard solution that has been widely investigated to implement hiding techniques. The possibility of employing hundreds of operating frequencies allows to effectively deform the side-channel trace at the point of preventing the correct pinpointing of the CP and, consequently, invalidating the possibility of attacking the target device even if the attacker can deploy an effective side-channel attack.
@@
  • We evaluate the effectiveness of our solution across various cryptographic primitives. Furthermore, we perform successful side-channel attacks after CP localization and alignment to validate the quality of the results and compare our solution against state-of-the-art techniques.
  • To facilitate further research and ensure reproducibility, we have released Hound as an open-source tool along with a dataset of relevant side-channel traces. The tool and dataset are available on GitHub 111 https://github.com/hardware-fab/Hound .
The rest of the paper is organized into four sections. Section II discusses the academic and commercial tools to locate the CPs within side-channel traces. Section III presents the proposed Hound approach. Section IV details the experimental results. Finally, Section V presents the conclusions.
@@ II Related Works
Figure 2: Focus on the proposed Hound training pipeline, divided into Dataset Building and CNN Training. $Class_{0}$, $class_{1}$ and, $class_{2}$ contain CP start part windows, CP spare parts windows, and noise windows, respectively.
However, using architectural-level techniques to morph the power trace represents an easy-to-implement and effective countermeasure to deceive pattern-matching-based solutions. For instance, employing time-sharing multithreading on a single-core microcontroller [22] can obstruct an attacker’s ability to accurately locate CPs, with interrupt service routines further modifying the side-channel trace’s form. Consequently, more sophisticated methods for detecting CPs in the side-channel trace have been developed. Chiari et al. [19] developed a deep-learning approach for identifying deformed CPs through random delay, employing a Convolutional Neural Network (CNN) trained to differentiate CPs from the remaining portions of the power trace. Even though random delay-based hiding techniques generate some degree of CP deformation, they are limited in their ability to introduce significant variability in the power trace [20].
Despite these advancements, current state-of-the-art techniques struggle with highly deformed traces, particularly when the computing platform employs effective hiding countermeasures like DFS [23]. These countermeasures profoundly alter the timing of operations and continuously morph the side-channel signal, making it challenging to isolate CPs using traditional methods. To address this limitation, we propose a deep-learning-based technique for CP localization in power traces collected from platforms implementing DFS as a countermeasure.
@@ III Methodology
Hound - Figure 1 depicts the proposed Hound methodology consisting of a training and inference pipeline. Figure 1a reports the Training Pipeline, which aims at creating a classifier that can sort a window of the side-channel trace into three categories: start of a CP, spare part of a CP, or just noise (see Figure 2). This framework locates where the CPs begin in the side-channel trace without removing the obfuscation countermeasure. We assume that the attacker leverages an exact copy of the target device to collect a noise trace and some cipher traces for the training. The noise trace is obtained by running general-purpose applications that are not the CP. Such information is necessary to train a neural network to distinguish the beginning of the CPs from the execution of other applications. Each cipher trace is gathered while a single CP is running, where the attacker decides what data and keys to use. All measurements are collected with the obfuscation mechanism active since the attacker cannot turn it off. Starting with measured raw traces, the Dataset Building stage creates a database of windows from the noise and cipher traces, each being $N$ samples long. This database is then used to train a CNN classifier.
Figure 1b reports the Inference Pipeline. At inference time, the goal is to identify the CPs within a novel side-channel trace from the target device by employing the trained CNN. This pipeline is divided into three key stages: Sliding Window Classification, Screening, and Alignment. The inference pipeline takes in a single side-channel trace and segments it to mark the start of each CP. Initially, the Slicing stage processes a side-channel trace and generates a series of windows to classify with the trained CNN. Starting from the classified windows, representing a segmentation of the input trace, the Screening stage produces a sequence of time instants marking the start of each CP in the trace. Finally, the Alignment stage chunks the original side-channel trace based on these outputs, arranging the identified CPs accordingly.
The details related to the dataset creation and the CNN architecture are discussed in Section III-A, while the sliding window classification and the screening procedure are detailed in Section III-B.
@@ III-A Training Pipeline
Dataset Building - The first stage in the training pipeline is Dataset Building, which receives a collection of side-channel traces as input and generates a dataset suitable for training the CNN. The attacker uses a clone of the target computing platform to create a noise trace and a set of cipher traces. The noise trace captures the execution of various general-purpose applications, excluding the target CP. Each cipher trace is collected while executing a single CP, where the attacker controls the plaintext and secret key. Notably, the attacker cannot disable the obfuscation mechanism, so the DFS remains active during all trace collection. Considering the proposed threat model, the attacker can only execute a chosen application and measure the corresponding side channel on the clone device. However, the attacker can probe the clone device to detect the start of each CP. In contrast, probing is not required for the noise trace. Once trained, the CNN works on the target architecture that implements DFS without the need for probing. For each cipher trace $i$ of length $L_{i}$ samples, the starting $N$ samples are labeled as start of the CP (see $class_{0}$ in Figure 2). The remaining $L_{i}-N$ samples are equally split into consecutive windows of width $N$ and labeled as spare part of the CP (see $class_{1}$ in Figure 2). Moreover, we extract a random set of $N$-sample windows from the noise trace and we label each of them as noise (see $class_{2}$ in Figure 2). Although the methodology aims to identify only the beginning of each CP, the division into three classes allows CNN to maximize the distinction over the three cases.
Convolutional Neural Network - The architecture of the 1D CNN is adapted from the CNN proposed in [19]. The CNN takes a window of $N$ samples from the side-channel trace as input and outputs a classification score vector. The CNN’s structure is organized into six sequential components: a convolutional block, a pair of residual blocks, a global average pooling layer, a fully-connected block, and a softmax layer. Each convolutional block features a 1D convolutional layer, a batch normalization layer [24], and a ReLU activation function. The residual blocks [25], each containing two convolutional blocks, are augmented with shortcut connections to perform element-wise feature summation. All the convolutional layers implement a kernel size of 64, a stride of 1, and zero padding to maintain the sample count to $N$. The first convolutional layer and the one in the first residual block implement 16 filters, while the subsequent residual block doubles the filter count to 32. The global average pooling layer averages the obtained features over the temporal dimension $N$. The feature vector is then processed through two fully-connected layers with a ReLU activation function. The softmax layer completes the architecture, generating a $3$-dimensional vector with the classification scores for each class.
@@ III-B Inference Pipeline
The Inference Pipeline, detailed in Figure 3, is divided into Sliding Window Classification, Screening, and Alignment stages. Leveraging the trained CNN classifier, the pipeline locates the starting sample of each CP present in a new side-channel trace. Sliding Window Classification - The first stage in the inference pipeline is Sliding Window Classification. It takes a new side-channel trace, slices it into $N$-sample windows, and uses the trained CNN to output a classification score to label each window as start of the CP, spare part of the CP, or noise (see Sliding Window Classification in Figure 3). The Slicing stage implements a sliding window procedure to slice the side-channel trace in input. It takes three inputs, i.e., the side-channel trace, the size of the sliding window ($N$), and the stride ($s$), and outputs an ordered set of $N$-sample windows to feed the CNN.
@@
Screening - The Screening stage refines the segmentation output from the Sliding Window Classification stage to identify the beginning of the CPs in the side-channel trace. Algorithm 1 highlights the screening procedure. The screening algorithm (see Algorithm 1) takes the segmentation output from the Sliding Window Classification stage, denoted as $seg$, and some parameters, i.e., the stride $s$ used for the sliding window, the kernel size $k$ for the polish procedure (see line 10 of Algorithm 1), and the average CP length $avgCP$. In the end, it returns the list $starts$ of samples corresponding to the beginning of each CP instance in the processed side-channel trace.
The algorithm is an iterative process that involves three main steps: POLISH, EXTRACT, and REFINE. The core idea behind Algorithm 1 is to start with an aggressive polish and incrementally refine it by attempting to uncover new CPs. The polishing step (see line 10 of Algorithm 1) employs a majority filter across the whole segmentation. This filter traverses the input signal for a given kernel size $k$, replacing each point with the most frequently occurring value within a $k$-sized window. The extraction step (see line 14 of Algorithm 1) identifies the initial sample of each CP. It pinpoints the indices of the falling edge down to class 0, meaning those samples with a value of 0 where the preceding value is not 0. These samples are then scaled by $s$, i.e., the stride in the sliding window classification. Finally, the refining step (see line 23 of Algorithm 1) fine-tunes the algorithm’s parameters by reducing kernel size $k$ and generating a list of sub-segmentation for subsequent iterations. These sub-segments encompass those regions of the input segmentation for which multiple hidden CPs might still exist. The process begins by determining the minimum length of a CP $minCP$ (see line 25 of Algorithm 1), which is derived by comparing the input’s average CP length and the CP’s starting points identified so far. Next, the algorithm examines pairs of consecutive CP starting points (see line 28 of Algorithm 1). If the distance between them exceeds twice the $minCP$, it indicates the potential presence of additional CPs within that interval. Consequently, the corresponding sub-segmentation is queued for further processing.
@@ IV Experimental Evaluation
TABLE I: Parameters for each pipeline stage and metrics score over all the tested ciphers.
Cipher General-Purpose Inference parameters CNN hyperparameters Metrics
Applications avgCP k N s batch size lr dropout Hits Mean IoU Std IoU
AES 145k 150 10k 62 256 0.01 0.2 100% 97.01% 1.65%
@@
This section presents the experimental results of Hound in four parts. Section IV-A outlines the experimental setup. Experimental results targeting the CNN training and inference pipelines are discussed in Section IV-B and Section IV-C, respectively. A comprehensive example workflow, including a successful side-channel attack and a comparison with state-of-the-art proposals, is provided in Section IV-D.